OpenStudy (anonymous):

Please can You Help? A die is rolled three times. i-) Find the chance that three different faces appear. My approach using complement rule: Probability p(same face pops up is)= 1- P. You roll first one and record what you get. The probability of next (2) two rolls give you same face is...? My chances of rolling any given face is 1/6. It does not matter what you get first role the next two rows have to the same as the first one role. Therefore= 1-(1/6)^2)= 1-(1/6^2)=

OpenStudy (anonymous):

i think the approach is right, but the answer may be wrong

OpenStudy (perl):

for the first die there are 6 possibilities, then for the second die there will be 5, then the third die there will be 4. since we are doing probability, the denominator will be 6^3 ( 6x5x4)/(6x6x6)

OpenStudy (anonymous):

oh maybe not, maybe the answer is right too roll one die. you get something roll another, the probability it is different it \(\frac{5}{6}\) then roll a third, the probability it is different from the first two is \(\frac{4}{6}\)

OpenStudy (anonymous):

multiply and get what @perl got, namely \[\frac{5\times 4}{6\times 6}\]

OpenStudy (perl):

using the complement would be more complicated

OpenStudy (anonymous):

Show me the multiple rule?